OpenStudy (anonymous):

Rationalize the Denominator. 9/square root of 2 and 4/square root of 6

OpenStudy (anonymous):

"and" means two different problems?

OpenStudy (anonymous):

\[\frac{9}{\sqrt{2}}\times \frac{\sqrt{2}}{\sqrt{2}}=\frac{9\sqrt{2}}{4}\]

OpenStudy (anonymous):

second one is similar

OpenStudy (anonymous):

what did you get for the second one

OpenStudy (anonymous):

4 square root of 6/6?

jhonyy9 (jhonyy9):

hello mr master radical 2 multiplied with radical 2 not equal 4 this just,only 2 ATTENTION !!!

jhonyy9 (jhonyy9):

hence the first will be 9radical2/2 and second 4radical 6/6 after you simplified with 2 you get 2radical6/3